OpenStudy (anonymous):

Solve each system by the addition method. 6x - 10y = 2 2x + 5y = 9 Put your answer in ordered pair (x, y) form.

OpenStudy (anonymous):

Before you add both systems together you should multiply the second equation by -3 to eliminate x. When you do that you should have : 6x -10y = 2 -6x - 15y = -27

OpenStudy (anonymous):

Now when you add this is what you get: -25y = -25 Therefore y = 1

OpenStudy (anonymous):

multiply the second equation by 2 is better

OpenStudy (anonymous):

If y = 1 we will substitute this value into either original equation 1 or 2. So lets pick equation 1: 6x - 10y = 2 if y = 1 we have 6x - 10 = 2 6x = 12 x = 2 So your final solution in ordered pair: (2,1)

OpenStudy (anonymous):

After your result you can check your solution by plugging in your result for both x and y so if we take the first equation and plug in 2 for x and 1 for y we get: 12 - 10 = 2 2=2 Since 2 is indeed equal to 2 your solution is correct

OpenStudy (anonymous):

As mathematicalpro rightly pointed out another way of solving is if you decide to eliminate y first. In that case you should multiply equation (2) by 2 and then add them together.

OpenStudy (anonymous):

Thanks so much!